# Embed-Search-Align: DNA Sequence Alignment using Transformer models

### v3

@@ Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to successfully align every read. We address this open problem by framing it as an “Embed-Search-Align” task. In this framework, a novel encoder model DNA-ESA generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. DNA-ESA is $>97\%$ accurate when aligning $250$-length reads onto a human reference genome of 3 gigabases (single-haploid), far exceeds the performance of $6$ recent DNA-Transformer model baselines and shows task transfer across chromosomes and species.
1 Background
@@
The aim of this paper is to establish a foundation model tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case). Numerous DNA Transformer models (Ji et al., 2021; Zvyagin et al., 2022; Fishman et al., 2023; Dalla-Torre et al., 2023) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as, genomes) and natural languages with large vocabulary. In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. Existing foundation Transformer models generate sequence embeddings such that their pairwise distances correspond to class separation, thus sequences with very large edit distances end up with representations that are close by. There are however tasks such as Sequence Alignment where the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 2, current DNA Transformer models fail to perform Sequence Alignment task. This requires a new kind of foundation model which is able to transfer a precise and computationally expensive distance metric over LVLs to a representation space.
The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads. Given a reference sequence :=$\mathcal{R}:=\{b_{1},b_{2},\dots,b_{N}\}$ – for the single-haploid human genome (Nurk et al., 2022), $N\approx 3$ gigabases (gb) – the primary objective is to identify the most probable start- and end-positions within this reference for a short DNA read,
@@
which may contain mutations due to base insertions, deletions, and substitutions. Computational simulators have been developed to generate synthetic reads that have properties of real reads. These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (Huang et al., 2012).
@@
(a) A $20,000$-long nucleotide sequence from Chr. 1, 2, and 3 each is randomly sampled and broken up into $100$ consecutive reference fragments (||||) each of length $1000$ and stride $200$. For each fragment, we sample a read (–) of 100-250 consecutive base pairs. Each sequence is encoded using DNA-ESA and the embeddings are visualized using 2D UMAP (McInnes et al., 2018). The colors identify chromosomes and the position of the fragment is coded by its intensity.
(b) Fragment (||||) and read (–) sequences from five gene regions (listed above in the inset) are similarly projected into a shared latent space. Note that for UMAP projections (both subplots (a) and (b)), the consecutive fragments belonging to the same nucleotide sequence constitute an order-preserving 1D manifold. A successfully aligned read is observed to be co-located with its corresponding fragment in the embedding space while a read further away is observed to be misclassified.
@@
Figure 1: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$. For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space. Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA.
@@
Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic Smith-Waterman (SW) approach, which computes similarity between a read $r$ and reference $\mathcal{R}$, but is computationally intensive for large genomes. Key advancements include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (Li & Durbin, 2009), reducing search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (Li, 2018; Vasimuddin et al., 2019; Langmead et al., 2018).
@@
In this work we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction the introduction of models like the Transformers (Vaswani et al., 2017) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures, as evidenced in various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (Devlin et al., 2019; Reimers & Gurevych, 2019; Qiu et al., 2020). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (Dosovitskiy et al., 2020), auditory (Verma & Berger, 2021), and neurological domains (Wang et al., 2023). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
@@ 1.1 Transformer models: Written Language to DNA Sequence Alignment
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (Huffman, 1952; Manber & Myers, 1993) – to successfully parse and compress DNA sequences. Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, genes.
Within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (Ji et al., 2021; Zhou et al., 2023), Nucleotide Transformer (Dalla-Torre et al., 2023), GenSLM (Zvyagin et al., 2022), and GENA-LM (Fishman et al., 2023) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to genetic function. Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment? We delineate the associated challenges as follows:
Figure 2: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling. Their performance, measured by recall (top-$K$) over $40K$ reads of varying lengths across the human genome (3gb - single-haploid), is shown. Trendlines represent each baseline, with error bars (Clopper-Pearson Interval (Clopper & Pearson, 1934) @ 95%) in grey. The vertical line at $x=250$ marks a typical read length. Overall, these baselines show suboptimal performance. For more details, see Sec. 4.
@@
    Two-Stage Training: DNA-based Transformer models typically undergo pretraining via a Next Token/Masked Token Prediction framework, a method originally developed for natural language tasks. To form sequence-level representations, these models often employ pooling techniques that aggregate token-level features into a single feature vector. This approach, however, is sometimes critiqued for yielding suboptimal aggregate features  (Reimers & Gurevych, 2019).
@@ 2 Our Contributions
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space. (B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank. This provides efficient access to the entire reference sequence for each read alignment. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as prose summarization and paragraph classification (Gao et al., 2021; Chen et al., 2020a). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (Johnson et al., 2019) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
3 Methods
@@
where $q\in\mathbb{N}_{0}$ is a candidate reference starting position and $v^{*}$ is the optimal alignment score. Lower scores indicate better alignments. This optimization exhibits the following property:
[P1] Sharding for sequence alignment: for a read segment $r$ of length $Q$ and reference $\mathcal{R}$ of length $N$, the complexity of $\texttt{SA}(r,\mathcal{R})$ scales as $\mathcal{O}(Q)$ when $N\rightarrow Q$.
Using P1, we can simplify the optimization problem by breaking it into sub-tasks with significantly shorter reference sequences. Specifically:
@@
Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\in\mathcal{R}$), and $K$ is the number of these sub-tasks [^2] We denote the relative distance between an alignment and the optimal score at read length $Q$ as $d_{SW}=\frac{|mQ-v^{*}|}{|mQ|}$ where $m(=-2)$ is the match score while computing the SW distance.. This approximation is effective under the conditions:
@@
    Fragment $\mathcal{F}_{j}$ lengths are on the order of the read length ($r$), not the longer reference ($\mathcal{R}$);
@@
Conditions (1) and (2) imply that fragments should be short and numerous enough to cover the reference genome. Condition (3) restricts the number of retrieved reference fragments per read — that we deem to be most likely to contain $r$ — to a small value $K$. Analogous methods have shown efficacy in text-based Search-and-Retrieval tasks  (Peng et al., 2023; Dai et al., 2022) on Open-Domain Question-Answering, Ranking among other tasks. Subsequent sections describe a parallel framework for retrieving reference fragments given a read. The pipeline is shown in Figs. 3 and 4.
Figure 3: System Overview [A] - Training Encoder and Populating Vector Store: Reference genome fragments $\mathcal{F}_{i}$ and within them, randomly sampled pure reads $r_{i}$ (positive pairs) are numerically represented via shared encoder $h$. Encoder training follows a contrastive approach as per equation 5. After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.

### v4

@@ Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully. We address this open problem by framing it as an “Embed-Search-Align” task. In this framework, a novel encoder model DNA-ESA generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, DNA-ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. DNA-ESA is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem. DNA-ESA far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
keywords:
Transformers, DNA Sequence Alignment, Large Language Models, Vector Stores
1 Introduction
Equal Contribution
Sequence alignment is a central problem in the analysis of sequence data. Many DNA sequencers generate short reads that are only a couple of hundred bases long. In order to interpret this data, a typical first step is to align the reads to a genome. Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long. Moreover, most experiments will generate millions of reads. Therefore, the resulting task of aligning reads to genomes is computationally expensive, and many decades of work have led to optimized approaches that can perform these tasks with great efficiency. However, many limitations remain with existing genome alignment methods. One limitation is that a reference genome usually is an approximation of the sequence from a single individual, and does not capture the variability of genomes across populations. Another, is that sequence alignment algorithms may not correctly model sequencing errors, and thus may introduce mistakes when trying to score the best match for a sequence against a genome. We asked whether some of these limitations could potentially be overcome if we utilize a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts. The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
@@ 1 Introduction
We establish a foundation model foundation tailored for DNA sequences, where the vocabulary consists of only a few symbols ($\{A,T,G,C\}$ in this case). Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Vocabulary Languages (LVL) (such as genomes) and natural languages with large vocabulary. In LVLs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. Existing foundation Transformer models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation. As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by. There are, however, tasks such as Sequence Alignment where the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective. This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LVLs to a continuous representation space.
The simplest sequence alignment task applies to single-end [^1] A DNA fragment is ligated to an adapter and then sequenced from one end only. reads, where a sequencer generates a read of length $Q$
$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
where $\tilde{b}_{i}\in\{A,T,G,C\}$. In practice, these reads come from individual genomes that do not necessarily match the reference and may contain mutations due to base insertions, deletions, and substitutions. Thus, it is assumed that this read is a noisy substring taken from a reference genome sequence $\mathcal{R}:=(b_{1},b_{2},\dots,b_{N}),\,b_{i}\in\{A,T,G,C\}$ and $N\gg Q$; for example, for the single-haploid human genome (chm_consort, ), $N\approx 3$ gigabases (gb), and $Q\approx 250$ though reads of much longer lengths are becoming increasingly affordable and accurate giab . The alignment task is to find a substring in $\mathcal{R}$
@@
such that the edit distance $d(r,\tilde{r})$ – for example the Smith-Waterman (SW) distance – is minimized. The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
Figure 1: Alignment Recall of Transformer-DNA Baselines by Read Length: Existing Transformer-DNA models were adapted for sequence alignment using mean-/max-pooling. Their performance, measured by recall (top-$K$) over $40K$ reads of varying lengths across the human genome (3gb - single-haploid), is shown. Trendlines represent each baseline, with error bars (Clopper-Pearson Interval (cps, ) @ 95%) in grey. The vertical line at $x=250$ marks a typical read length. Overall, these baselines show suboptimal performance. For more details, see Sec. 3.
@@
Computational simulators have been developed to generate synthetic reads that have properties of real reads. These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (huang_art_2012, ).
@@
((a)) Three $20,000$-long nucleotide sequences were selected, one each from Chr. 1, 2, and 3. Each sequence is broken up into $100$ consecutive reference fragments, each of length $1000$. Consecutive fragments have an overlap of $800$ bases (stride $200$). Additionally, reads of length $Q\sim\mathcal{U}[100,250]$ are sampled from within each fragment. The colors correspond to the respective chromosomes and the position of a particular fragment along the 20K-nucleotide sequence is coded by its color intensity.
((b)) Five gene regions (listed above in the inset) of different sizes are selected from the reference genome. Each gene is broken up into consecutive fragments of length $1000$ with a stride of $200$ (similar to subfigure (a)). Additionally, reads of length $Q\sim\mathcal{U}[100,250]$ are sampled from within each fragment. The colors correspond to genes and the position of a particular fragment along a specific gene is coded by its color intensity.
@@
Figure 2: Illustrating DNA-ESA’s Preservation of Sequence Locality in Embedding Space: In DNA-ESA, the reference genome $\mathcal{R}$ is divided into fragments $\mathcal{F}_{i}$, each represented by an embedding $h(\mathcal{F}_{i})$. For effective sequence alignment, specific structures are expected in the embedding space: (1) Overlapping fragments $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ should have proximate embeddings; (2) Consecutive fragments forming a long sequence should correspond to a distinct manifold in the embedding space. Subfigures (a) and (b) display this emergent geometry, as visualized using DNA-ESA. In both subfigures, the fragment and read representations are jointly visualized in a 2D low-dimensional space constructed by applying UMAP umap to the original embeddings. Vertical marker lines (||||) refer to fragments, “–” markers correspond to reads. We observe that the consecutive fragments belonging to the same nucleotide (subfigure (a)) or gene (subfigure (b)) sequence constitute an order-preserving 1D manifold. Additionally, almost all of the reads are close to their corresponding fragments, making this space viable for the task of Sequence Alignment: by searching for fragments in the neighborhood of an external read represented in the embedding space, one is likely to retrieve a fragment most responsible for the read.
@@
Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic SW approach, which computes the similarity between a read $r$ and reference $\mathcal{R}$; using the SW distance on its own, however, is computationally intensive for large genomes. Some key advances include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (bw_transform, ), reducing the effective search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (minimap, ; bwamem2, ; bowtie, ).
@@
In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020, ), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
@@ 1.1 Transformer models: Written Language to DNA Sequence Alignment
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Classical DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences. Furthermore, just as written language contains repeated subsequences (words, phrases) to represent real-world objects, DNA sequences similarly possess repeating “words” and groupings of such words into a “sentence” representing, for example, subsets of genes.
Within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function. Given these advances, a natural question arises: Can these Transformer architectures be readily applied to the task of Sequence Alignment? We delineate the associated challenges as follows:
@@
    Two-Stage Training: DNA-based Transformer models typically undergo pretraining via a Next Token/Masked Token Prediction framework, a method originally developed for natural language tasks. To form sequence-level representations, these models often employ pooling techniques that aggregate token-level features into a single feature vector. This approach, however, is sometimes critiqued for yielding suboptimal aggregate features  (reimers_sentence-bert_2019, ).
@@ 2 Our Contributions
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder DNA-ESA, trained through self-supervision, to map DNA reads to relevant fragments in a reference sequence within a shared embedding space. (B) We leverage a specialized data structure, termed a DNA vector store, as a memory bank [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/. This provides efficient access to the entire reference sequence for each read alignment. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
2.1 Task Definition
@@
where $q\in\{1,2,3,\dots\}$ is a candidate reference starting position and $v^{*}$ is the optimal alignment score. Lower scores indicate better alignments. This optimization exhibits the following property:
Sharding for sequence alignment: For a read segment $r$ of length $Q$ and reference $\mathcal{R}$ of length $N$, the complexity of $\texttt{SA}(r,\mathcal{R})$ scales as $\mathcal{O}(NQ)$, which is expensive.
Ideally, we would like to use $\mathcal{R}$ to get a set of fragments $\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{K}\}$ that are a subset of the original reference. Then:
@@
Here, each $\mathcal{F}_{j}$ is a fragment of $\mathcal{R}$ (i.e., $\mathcal{F}_{j}\subset\mathcal{R}$), and $K$ is the number of these sub-tasks. This approximation is effective under the conditions:
@@
    Fragment $\mathcal{F}_{j}$ lengths are on the order of the read length ($Q$), and not the length of the longer reference $\mathcal{R}$;
@@
Conditions (1) and (2) imply that fragments should be short and numerous enough to cover the reference genome. Condition (3) restricts the number of retrieved reference fragments per read — that we deem to be most likely to contain $r$ — to a small value $K$. Analogous methods have shown efficacy in text-based Search-and-Retrieval tasks  (peng2023, ; dai2022, ) on Open-Domain Question-Answering, Ranking among other tasks. Subsequent sections describe a parallel framework for retrieving reference fragments given a read. The pipeline is shown in Figs. 3 and 4.
Figure 3: System Overview [A] - Training Encoder and Populating Vector Store: Reference genome fragments $\mathcal{F}_{i}$ and within them, randomly sampled pure reads $r_{i}$ (positive pairs) are numerically represented via shared encoder $h$. Encoder training follows a contrastive approach as per (6). After training, the genome is segmented into overlapping fragments, encoded, and uploaded into the vector store.